08
Mar
2019
11:40

We apologise to customers for the brief outage caused by one of our suppliers between London and Maidenhead.

Everything is back to normal now.

Posted by Paul Malkowski

07
Mar
2019
03:33

At approximately 23:30 on 06/03/19 we lost connectivity between London and Maidenhead, affecting many services. The link is comprised of 2 connections from different suppliers. Unfortunately both of these connections were unavailable for a period, for reasons that are not currently fully understood.

We are working with both link suppliers and will post a fuller update when further information is available.

All services are now restored – if you should see anything unexpected please contact support on 020 7517 4999 or support@watchfront.co.uk

Aplogies for the inconvenience caused.

Posted by Kevin Hones

06
Mar
2019
08:46

Apologies to all our customer who have lines with us and experienced packet loss. There was an issue with one of our connections in London which has now been fixed.

Posted by Yolanda Zyczynska

04
Mar
2019
14:10

Apologies to all Leased Line and DSL customers who suffered a brief outage around 13:45 today.

This was caused by an interruption between London and Maidenhead by our suppliers having a network glitch.

Posted by Paul Malkowski

03
Oct
2018
12:05

Apologies to our customers who suffered a brief outage this morning.

Unfortunately, whilst Kev as conducting some housekeeping work in Blue Square 1 he accidentality rebooted the Peregrine switch. As this was a totally accidental occurrence we do not anticipate there to be a repeat.

We apologise for any inconvenience this may have caused.

Posted by Yolanda Zyczynska

02
Sep
2018
23:14

Update 07/09/18: The faulty unit was replaced yesterday and has been functioning normally sine.

One half of switch cluster ‘willem’ in Maidenhead is currently unresponsive. Any customers with single-homed CoLo connected to the faulty unit will currently be suffering loss of connectivity.
We will investigate and repair as soon as possible and will publish an update in the morning.
In the meantime, apologies for the inconvenience.

Posted by Kevin Hones

29
May
2018
11:04

Hello,

We suffered a DDoS attack on our network in the early hours of Monday the 28th of May.

One of our upstream providers noticed this and null routed the attack.

All services were back to normal in under two hours.

We apologise for any inconvenience this may have caused.

Posted by Paul Malkowski

10
Apr
2018
11:54

Update: All affected lines have now resumed normal service.

Customers who have a TalkTalk connected service may have experienced a glitch in service for a minute or so at around 11:00 this morning, services are slowly returning to normal to all affected lines.

The cause of the glitch is unknown at this time but is being investigated, as soon as we have more information from TalkTalk we will post an update.

We apologise for any inconvenience caused.

Posted by Yolanda Zyczynska

25
Jan
2018
11:21

Customers who have a TalkTalk connected service may have experienced a glitch in service for a minute or so at around 11:00 this morning, normal service is now running.

The cause of the glitch is unknown at this time, as soon as we have more information from TalkTalk we will post an update.

We apologise for any inconvenience caused.

Posted by Yolanda Zyczynska

27
Nov
2017
08:53

This morning TalkTalk are having a big outage which is in turn is making any EoFTTC or Leased Line from them unavailable.

We apologise for this problem and are assured TalkTalk are working to fix the issue as soon as possible,

We will update you here once we know more.

Update 09:13 27/11/2017

All lines have come back now and the problem is gone.

Apologies once more for this issue.

Posted by Paul Malkowski

17
Nov
2017
12:19

We apologise to our customers with BT lines who have been affected by the outage this afternoon. Our wholesale supplier for BT connected lines is investigating the cause of the issue.

Apologies to those customers affected.

Posted by Yolanda Zyczynska

29
Aug
2017
08:44

We are observing packet loss affecting cartain data paths in our AMS datacentre, which is being investigated.

Posted by Kevin Hones

Current Status

    GoodNo current issues

Status Archive